OpenStudy (anonymous):

Evaluate and, if possible, simplify. -x - 5/15 for x = -5

OpenStudy (anonymous):

0

OpenStudy (anonymous):

-(-5)-5/15 = 5-5/15 =0/15 =0

OpenStudy (anonymous):

plz show work

OpenStudy (anonymous):

^ just did.

OpenStudy (anonymous):

is it -x - (5/15) or (-x - 5)/15 Please use brackets

OpenStudy (radar):

Good point cuddlepony, the problem as presented is ambiguous.

OpenStudy (radar):

Answer might be 0 or maybe 4 2/3